OpenStudy (anonymous):

Help with the following ecuation: |2|x-a|-bx|=|2b-|x-a|| | | means absolite value Where b<0 and -2

OpenStudy (anonymous):

wow this is a mess

OpenStudy (anonymous):

what are you trying to find though

OpenStudy (anonymous):

We have to solve for x. So, seems possible solve for x in terms of b and a. And then we are suposed to define intervals for x, given the conditions for b and a. I think.

OpenStudy (anonymous):

ok on the right hand side if you take that part of the equation out of the absolute values it becomes|2|x-a|-bx| =2b+|x-a| correct?

OpenStudy (anonymous):

now we do that for the left side too so it becomes 2|x-a|+bx=2b+|x-a|

OpenStudy (anonymous):

Should't be 2|x-a|+bx=2b+|x-a| ?

OpenStudy (anonymous):

2|x-a|-bx=2b+|x-a| *

OpenStudy (anonymous):

no because the -bx was in one of the absolute values once anything comes out of an absolute value it instantly becomes positive.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

look at your original problem if you want to make sure

OpenStudy (anonymous):

now we still have 2 more absolute values to take care of. so can you tell me what the problem would look like if you took those out.

OpenStudy (anonymous):

2|x-a|+bx=2b+|x-a| |x-a|=2b-bx That means x-a=2b-bx x(1+b)=a+2b x= (a+2b)/(1+b) or a-x=2b-bx x(b-1)=2b-a x=(2b-a)/(b-1) I think its that

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

your second answer is close

OpenStudy (anonymous):

this is what i got 2|x-a|+bx=2b+|x-a| 2(x+a)+bx=2b+x+a 2x+2a+bx=2b+x+a 2x+bx-x=2b-a x(2+b-1)=2b-a \[x=\frac{ 2b-a }{ 2+b-1 }\]

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

hmm, i don't know, i thought that since the definition of absolute value is |x|={ x if x>0 {-x if x ≤0 we should find more than one solution for this ecuation. At least to define x in terms of a and b. And i dont know why you only take one part of the definition.

OpenStudy (anonymous):

let me look at it again i might have overlooked something

OpenStudy (anonymous):

hmmm crap i might have read it wrong

OpenStudy (anonymous):

i cant tell what exactly went wrong sorry

OpenStudy (anonymous):

from my result b cannot = -1 but thats all i can get from it

OpenStudy (anonymous):

sorry that i couldnt be much help

OpenStudy (anonymous):

There are 4 candidate solutions that will become apparent. Only 1 of the 4 will end up working, but I will demonstrate that. The solution that does work I will show works for the given interval and I will show that by proving the general case: all "a" and "b" for the given intervals for each. The other 3 can be shown not to work by simply selecting convenient values for "a" and "b". So, by eliminating the outer || : 2|x - a| - bx = 2b - |x - a| or -2b + |x - a| The first (P) of these will split into cases 1 & 2, the second (Q) will split into cases 3 & 4. I'll show eventually that only case 1 will be a solution. Taking P: |x - a| = bx - 2b x - a = bx - 2b (case #1) or -bx + 2b (case #2) case #1 : x = (2b - a)/(b - 1) case #2 : x = (2b + a)/(b + 1) Taking Q: 3|x - a| = 2b + bx 3x - 3a = 2b + bx (case #3) or 3x - 3a = -2b - bx (case #4) case #3 : x = (2b + 3a)/(3 - b) case #4 : x = (3a - 2b)/(3 + b) Now, proving a case DOESN'T work is easier, so I'll continue by selecting suitable, convenient values for "a" and "b" for cases #2, #3, and #4. You can plug them in and verify that the equality will not hold. This will be sufficient for these cases because the equality is supposed to hold over the range of all values for "a" and "b". So for each, we only need one counter-example: For case #2, select b = -2 and a = 0 For case #3, select b = -1 and a = 0 For case #4, select b = -2 and a = 0 Case #1 is the only one that works, but we need to show that for the general case: substituting back x = (2b - a)/(b - 1) into the original equation, the left side becomes:\[\left| 2\left| \frac{ 2b - a }{ b - 1 } - a \right| - b \left( \frac{ 2b - a }{ b - 1 } \right) \right| = \left| 2\left| \frac{ 2b - ab }{ b - 1 } \right| - \frac{ b }{ b - 1 }(2b - a) \right|\]\[= \left| 2\left( \frac{ b }{ b - 1 } \right)\left| 2 - a \right| - \left( \frac{ b }{ b - 1 } \right)(2b - a) \right|\]We can pull that b/(b-1) out of the || because it is always positive\[= \left( \frac{ b }{ b - 1 } \right)\left| (4 - 2a) + (a - 2b) \right|\]Likewise, "2 - a" is never negative.\[= \left( \frac{ b }{ b - 1 } \right)(4 - a - 2b)\]The right side will become the same as the immediate above:\[\left| 2b - \left| \frac{ 2b - a }{ b - 1 } - a \right| \right| = \left| 2b - \left| \frac{ 2b - ab }{ b - 1 } \right| \right|\]\[= \left| 2b - \left( \frac{ b }{ b - 1 } \right)\left| 2 - a \right| \right| = \left| 2b - \left( \frac{ b }{ b - 1 } \right)(2 - a) \right|\]\[= \left| 2b + \left( \frac{ b }{ b - 1 } \right)(a - 2) \right| = \left| \frac{ 2b ^{2} - 2b + ab - 2b }{ b - 1 } \right|\]\[= \left( \frac{ b }{ b - 1 } \right)\left| 2b - 4 + a \right| = \left( \frac{ b }{ b - 1 } \right)\left| -(4 - a - 2b) \right|\]\[= \left( \frac{ b }{ b - 1 } \right)(4 - a - 2b)\]

OpenStudy (anonymous):

I just saw that "taking P" is "taking Q" and vice versa, I just copied it out a little out of order, but I'm sure you'll see that. @juarismi

OpenStudy (anonymous):

Great, many thanks. Now i see it!!. But just one last thing. if the only solution is that works is x =(2b - a)/(b - 1) and b<0 and -2<a≤2 That means that x will always be ≤2 ¿?

OpenStudy (anonymous):

Ok, thank you so much.

OpenStudy (anonymous):

uw! Great working with you and I hope this helps a lot!

OpenStudy (anonymous):

@juarismi , I think you might be right that x <= 2

OpenStudy (anonymous):

Me too. I can't find a case where x>2.

OpenStudy (anonymous):

Yes, you are right because if we take "a" as large as possible (a = 2) and "b" as large as possible (b approaching 0), the largest we get is not even as large as: [2(0) - (2)] / (0 - 1) = 2

OpenStudy (anonymous):

Fine. Thanks

OpenStudy (anonymous):

I have a better answer than what I gave: Maximize "a" to be "2" y = (2b - 2) / (b - 1) = 2 So, when "a" = 2, all "b" give x = 2

OpenStudy (anonymous):

Do you agree? That way, because the denominator is always negative, we want to maximize the absolute value of the numerator by making it "more negative" and that is by making "a" as large as possible so we can subtract the max. @juarismi

OpenStudy (anonymous):

I agree. Absolutely. So i think the answer can be write as: is x =(2b - a)/(b - 1) and x is <=2.

OpenStudy (anonymous):

Yes. I think we nailed this one!